When plugged into a standard U.S. outlet (120 V), a motor draws 0.750 A of current. If this motor is 35% efficient, how long does it take for this motor to lift an object with a mass of 2.90 kg to a height of 3.00 m?

THE ANSWER SHOULD BE 2.71 SECONDS. PLEASE EXPLAIN STEP BY STEP AND WHY IF YOU COULD.
Answer #1

given that

potential difference V = 120 V

current i = 0.75 A

mass m = 2.9 kg

height h = 3 m

from the relation power p = V*i

p = 120*0.75 = 90 W

but the efficiency is 35%

so power p = 0.35*90 = 31.5 W

power p = energy/time

p = mgh/t

time t = 2.9*9.8*3/31.5

t = 2.71 sec